### v2

# A Universal Scheme for Dynamic Partitioned Shortest Path Index

Abstract.
Shortest path (SP) computation is the fundamental operation in various networks such as urban networks, logistic networks, communication networks, social networks, etc. With the development of technology and societal expansions, those networks tend to be massive. This, in turn, causes deteriorated performance of SP computation, and graph partitioning is commonly leveraged to scale up the SP algorithms. However, the partitioned shortest path (PSP) index has never been systematically investigated and theoretically analyzed, and there is a lack of experimental comparison among different PSP indexes. Moreover, few studies have explored PSP index maintenance in dynamic networks. Therefore, in this paper, we systematically analyze the dynamic PSP index by proposing a universal scheme for it. Specifically, we first propose two novel partitioned shortest path strategies (No-boundary and Post-boundary strategies) to improve the performance of PSP indexes and design the corresponding index maintenance approaches to deal with dynamic scenarios. Then we categorize the partition methods from the perspective of partition structure to facilitate the selection of partition methods in the PSP index. Furthermore, we propose a universal scheme for designing the PSP index by coupling its three dimensions (i.e. PSP strategy, partition structure, and SP algorithm). Based on this scheme, we propose five new PSP indexes with prominent performance in either query or update efficiency. Lastly, extensive experiments are implemented to demonstrate the effectiveness of the proposed PSP scheme, with valuable guidance provided on the PSP index design.
PVLDB Reference Format:
PVLDB, (): , .
doi: This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. , No.  ISSN 2150-8097.
doi:
footnotetext: *Both authors contributed equally to this paper. Lei Li is the corresponding author.
1. Introduction
Shortest Path (SP) query in dynamic networks is an essential building block for various applications affecting daily lives. Given an origin-destination pair, it returns the minimum traveling time in a road network (Bast et al., 2016), the fastest connection in the web graph (Boccaletti et al., 2006), or the most intimate relationships in the social network (Kolaczyk et al., 2009; Opsahl et al., 2010; Yao et al., 2013). With the development of urban traffic systems and the evolvement of online interactions platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the state-of-the-art (Li et al., 2020a, 2019b; Zhang et al., 2021a, c, b, 2023a) with either heavy memory or expensive search overheads. For example, 2-Hop Labeling (Cohen et al., 2003) has the fastest query efficiency but requires $O(nm^{1/2})$ space (for a graph with $n$ vertices and $m$ edges) to store the index. Complementary, the search-based methods like Dijkstra’s are inefficient on query processing on large networks. Moreover, because the networks are dynamic in nature with evolving structures or edge weights, extra efforts are required to handle the updates for SP indexes.
Graph partitioning is often used to improve the scalability of SP algorithms, enabling more complicated problems like time-dependent (Li et al., 2019c, 2020b) and constraint path computation (Wang et al., 2016; Liu et al., 2021, 2022a) to be efficiently handled in large networks. It decomposes a large network into several smaller ones such that the construction time can be improved by parallelization, and the index sizes and update workloads can be reduced. As shown in Figure 1, we can compare them briefly in terms of index construction time, storage space, query time, and update time. On one extreme are the direct search algorithms (Dijkstra, 1959; Hart et al., 1968; Li et al., 2020c; Zhang et al., 2020) that require no index and can work in dynamic environments directly but are slow for query answering. Then the partitioned search adds various information to guide and reduce search space (Chondrogiannis and Gamper, 2016; Möhring et al., 2007; Bauer and Delling, 2010) to improve query processing. On the other extreme, Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-hop Labeling (HL) (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance. Their partitioned versions (Delling et al., 2017; Li et al., 2020a; Li et al., 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2022b, 2022a) are faster to construct with smaller sizes but longer query times. As these works have better performance than their unpartitioned versions, they create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches; 2) Different PSP indexes are irrelevant and very different from each other; 3) Applying graph partitioning is always better.
Figure 1. Performance Comparison of SP Algorithms
Motivations. These misconceptions are not always true. Although the partitioned shortest path (PSP) indexes have been widely applied in the past decade, to the best of our knowledge, they have not been studied systematically. Specifically, the existing solutions typically pick one partition method without discriminating their characteristics and figure out a way to make it work, and then claim benefits/superiority. Consequently, there lacks a generalized scheme to organize and compare the PSP indexes insightfully and fairly. Towards that, we postulate that both theoretical and experimental systematic study is needed for PSP indexes, so that a suitable index could be constructed for different scenarios.
Challenges. However, it is non-trivial to design such a scheme. Firstly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority. However, since no effort has been put into the abstract PSP index itself, it is unclear what determines their performance, when to use particular indexes, and how to design a suitable new PSP index for a new scenario. Therefore, we study the PSP index systematically in this work and identify three dimensions throughout the rich literature. Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance. Therefore, we dive into the partitioned shortest path (PSP) strategy, which determines how to coordinate multiple partitions for correct query answering, and summarizes the traditional PSP strategy. Additionally, we propose two novel PSP strategies together with their index update methods to improve the index performance. Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further. Thirdly, the graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, etc.), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path-oriented applications. Therefore, we propose a path-oriented partitioning classification from the perspective of partition structure for easier partition method selection. Lastly, after elaborating on each dimension of PSP index, we propose our universal PSP index scheme by entangling these dimensions. Based on this scheme, we propose five new PSP indexes to achieve better efficiency in either query or update operations by selecting the specific strategy or method on each dimension as per system requirements.
Contributions. Our contributions are listed as follows:
  • $\bullet$
    We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: PSP strategy, partition structure, and SP algorithm, such that the PSP index can be analyzed and designed systematically. (Section 2)
  • $\bullet$
    We propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve the index construction and maintenance efficiency over the traditional PSP strategy. We also provide non-trivial correctness analysis for their index construction, query processing, and index maintenance. A novel Pruning-based overlay graph optimization is also designed to prune the unnecessary computation. (Section 3)
  • $\bullet$
    We propose a new path-oriented partitioning classification from the perspective of partition structure for easier partition method selection in designing the PSP index. (Section 4)
  • $\bullet$
    Based on the proposed PSP index scheme, we put forward a PSP index generator to create new PSP structures. Among them, we identify and design five new PSP indexes suitable for specific network structures and scenarios. (Section 5)
  • $\bullet$
    We implement comprehensive experimental studies to explore the performance of PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios. (Section 6)
2. Preliminary and PSP Dimensions
In this paper, we focus on the dynamic weighted network $G(V,E,W)$ with $V$ denoting the vertex set, $E$ denoting the edge set and $W:E\rightarrow\mathbf{R}^{+}$ assigning a non-negative weight to each edge – i.e. $w(u,v)\in W,\forall(u,v)\in E$ – which can increase or decrease in an ad-hoc manner. We denote the number of vertices and edges as $n=|V|$ and $m=|E|$. For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$, u’s commonly called adjacent to $v$. We associate each vertex $v\in V$ with order $r(v)$ indicating its importance in $G$. In this paper, we use boundary-first ordering (Liu et al., 2022a) to generate the order for the PSP indexes. For non-partitioned SP indexes, we leverage degree to decide the order of small-world networks while the minimum degree elimination-based ordering (Berry et al., 2003) for the road networks. A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$. Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e. $p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$. In the following, we introduce the PSP index dimensions and summarize each of them.
2.1. Dimensions of PSP Index
Based on a comprehensive literature review and analysis of the PSP index, we identify three critical dimensions of the PSP index: 1) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes. However, the existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study; 2) PSP Strategy, which provides a procedure to organize the index construction for multiple subgraphs and coordinate those indexes such that the query correctness is guaranteed. Only straightforward solutions exist currently as will be discussed in Section 2.3. However, this part is crucial because it determines the complexity of both the query and update; 3) SP Algorithm, whose influence was thoroughly studied in our previous work (Zhang et al., 2021a). With these three dimensions, we can form a space to position the existing methods into it. Given a massive dynamic network, these dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query/update efficiency) can be satisfied. Next, we discuss and review these three dimensions.
2.2. Graph Partitioning
Definition 0 (Graph Partitioning).
A graph $G$ is decomposed into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with $V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$. denoted as $E_{inter}$ and $E_{intra}$.
We now discuss and categorize the vertex and edge in graph partitioning. $\forall v\in G_{i}$, we say $v$ is a boundary vertex if there exists a neighbor of $v$ in the another subgraph, that is $\exists u\in N(v),u\in G_{j}(i\neq j)$. Otherwise, $v$ is an inner vertex. We represent the boundary vertex set of $G_{i}$ as $B_{i}$ and that of $G$ as $B=\bigcup B_{i}$. For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs, i.e., $u\in B_{i},v\in B_{j},i\neq j$. Otherwise, it is an intra-edge. The corresponding edge sets are denoted as $E_{inter}$ and $E_{intra}$, respectively. For example in Figure 2, the example graph is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundaries ($B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices. Edges outside partitions like $(v_{3},v_{5})$ are inter-edges, and $(v_{3},v_{10})$ is an intra-edge. Besides, the boundary vertex plays an important role in the shortest path computation with its cut property: given a shortest path $p(s,t)$ with its endpoints in two different subgraphs $s\in G_{i},t\notin G_{i}$, there is at least one vertex $v$ which is in the boundary vertex set of $G_{i}$ on the path, that is $\exists v\in p(s,t),v\in B_{i}$. Therefore, boundary vertices would determine the correctness and complexity of partitioned shortest path computation (discussed in the following sections). Various partition methods are involved in the existing PSP methods such as METIS (Karypis and Kumar, 1998) in G-Tree (Zhong et al., 2013) and SHARC (Bauer and Delling, 2010) ; PUNCH (Delling et al., 2011) in COLA (Wang et al., 2016), FHL (Liu et al., 2021, 2022b), FHL-Cube (Liu et al., 2022a) and THop (Li et al., 2019c, 2020b) ; Tree Decomposition-based Partition (Ouyang et al., 2018) in Core-Tree (Li et al., 2020a, 2019b) ; multi-level partitioning (Jung and Pramanik, 1996) in G-tree (Zhong et al., 2013, 2015), SPAH (Jung and Pramanik, 1996), CPR (Delling et al., 2017) and (Holzer et al., 2009; Chondrogiannis and Gamper, 2016, 2014). But, their selections are performed in an ad-hoc manner or consider the boundary number, partition balance, and spatial awareness, rather than the SP computation itself.
2.3. Traditional PSP Strategy
The PSP index $L$ typically consists of two components: the partition indexes $\{L_{i}\}$ for each subgraph (partition) $G_{i}$, and the overlay index $\tilde{L}$ for the overlay graph which is composed of the boundary vertices of all partitions, i.e., $L=\{L_{i}\}\cup\tilde{L}$. Regardless of the specific SP algorithm, almost all the existing PSP indexes (Jung and Pramanik, 2002; Delling et al., 2017; Holzer et al., 2009; Chondrogiannis and Gamper, 2014, 2016; Zhong et al., 2013; Li et al., 2019a; Zhong et al., 2015; Ta et al., 2017; Wang et al., 2019, 2016; Liu et al., 2021, 2022b, 2022a; Li et al., 2019c, 2020b; Li et al., 2020a; Zheng et al., 2022; Wang et al., 2021; Farhan et al., 2022; Gubichev et al., 2010; Qiao et al., 2012) construct the PSP index and process the queries under the following procedure.
Figure 2. Example Graph $G$ and Overlay Graph $\tilde{G}$
Index Construction. It first leverages a partition method to divide the graph $G$ into multiple subgraphs $\{G_{i}\}$ and then builds the PSP index $L$ by the following four steps. As illustrated in Figure 3, Step 1 precomputes the global distance between all boundary vertex pairs $(b_{i1},b_{i2}),(b_{i1},b_{i2}\in B_{i})$ for each partition $G_{i}$ and insert shortcuts $e(b_{i1},b_{i2})=d_{G}(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$. For instance, the global distance between boundary vertex pairs $(v_{3},v_{10})\in G_{1}$, $(v_{8},v_{9})\in G_{2}$, $(v_{5},v_{12})\in G_{3}$, $(v_{6},v_{11})\in G_{4}$ are calculated and inserted into their corresponding subgraphs to form new subgraphs $G^{\prime}_{1},G^{\prime}_{2},G^{\prime}_{3},G^{\prime}_{4}$, respectively; Step 2 constructs the SP index $L_{i}$ based on $G^{\prime}_{i}$. The index can be any type like CH (Geisberger et al., 2008), PLL (Akiba et al., 2013) or H2H (Ouyang et al., 2018), and the index construction procedure among multiple partitions can be parallelized as it takes each partition as input independently; Step 3 constructs the overlay graph $\tilde{G}$ based on the precomputed shortcuts in Step 1. Specifically, $\tilde{G}$ is composed of shortcuts between boundary vertex pairs $\{B_{i}\times B_{i}\}$ ($(v_{3},v_{10}),(v_{8},v_{9}),(v_{5},v_{12}),(v_{6},v_{11})$) and the inter-edge set $E_{inter}$ ($(v_{3},v_{5})$, $(v_{10},v_{12})$, $(v_{9},v_{10})$,$(v_{9},v_{12})$, $(v_{9},v_{11})$, $(v_{8},v_{11})$, $(v_{12},v_{11})$, $(v_{12},v_{6})$), that is $V_{\tilde{G}}=\{B_{i}\},E_{\tilde{G}}=\{B_{i}\times B_{i}\}\cup E_{inter}$; Step 4 constructs the SP index $\tilde{L}$ on $\tilde{G}$. Similar to $L_{i}$, $\tilde{L}$ can be any index type.
Figure 3. Traditional PSP Strategy Example
Note that the construction of $\tilde{L}$ (Step 3, Step 4) can be parallelized with $L_{i}$ (Step 2) since they are independent and both rely on Step 1. All pairs of boundaries are independent of each other so we only need $|B|$ times of Dijkstra’s with time complexity $O(nlogn+m)$. Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent to them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
Query Processing. By referring to the cut property of boundary vertex, the shortest path queries are divided into two categories and processed by utilizing $L$:
Case 1: Same-Partition i.e., $\forall s,t\in G_{i}$, $Q(s,t)=d_{L_{i}}(s,t)$;
Case 2: Cross-Partition, i.e., $\forall s\in G_{i},t\in G_{j}(i\neq j),Q(s,t)=$
$\begin{cases}d_{\tilde{L}}(s,t)&s,t\in B\\ \min\limits_{b_{q}\in B_{j}}\{d_{\tilde{L}}(s,b_{q})+d_{L_{j}}(b_{q},t)\}&s\in B,t\notin B\\ \min\limits_{b_{p}\in B_{i}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},t)\}&s\notin B,t\in B\\ \min\limits_{b_{p}\in B_{i},b_{q}\in B_{j}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},b_{q})+d_{L_{j}}(b_{q},t)\}&s\notin B,t\notin B\\ \end{cases}$
In summary, when $s$ and $t$ are in the same partition, we can use $L_{i}$ to answer $d_{G}(s,t)$; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The same-partition query complexity is $O_{q}(L_{i})$, where $O_{q}$ is the index’s query complexity. The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. The proof of the query processing can be found in the Appendix of our extended version (Zhang et al., 2023b).
2.4. Shortest Path Algorithms
We summarize the shortest path algorithms as follows: 1) Direct Search such as Dijkstra’s (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) searches the graph when the index cannot be constructed due to the large graph size (QbS (Wang et al., 2021), ParDist (Chondrogiannis and Gamper, 2016)) or complicated problem (COLA (Wang et al., 2016)). It takes no time in index update, but queries are slowest; 2) Contraction Hierarchy (CH) (Geisberger et al., 2008) is a widely used lightweight index that contracts the vertices in a pre-defined order and preserves the shortest path information by adding shortcuts among the contracted vertex’s neighbors. The search-based CH (Geisberger et al., 2008) has a small index size but takes longer to build and maintain (Zhang et al., 2021a; Geisberger et al., 2012) while concatenation-based CH (Ouyang et al., 2020; Wei et al., 2020) is much faster to build and maintain if the tree-width is small. CH’s query performance is generally around 10$\times$ faster than the direct search but much slower than the hub labelings. Surprisingly, no PSP index has used CH as underlying index; 3) Pruned Landmark Labeling (PLL). Although many hub labeling methods have been proposed in the past decade, only two are widely used. Built by either pruned search (Akiba et al., 2013) or propagation (Li et al., 2019b; Jin et al., 2020; Zhang et al., 2021b, 2023a), PLL is the only index that can work on the graph with large treewidth. Therefore, graphs with this property use it as the underlying index (e.g., QsB (Wang et al., 2021; Farhan et al., 2022) for its landmarks, Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) for its core); 4) Tree Decomposition (TD) (Ouyang et al., 2018; Chang et al., 2012; Wei, 2012; Chen et al., 2021; Zhang et al., 2021c). As another widely used hop labeling, it is much faster than PLL for graphs with smaller treewidth but cannot scale to large ones. For instance, CT (Li et al., 2020a; Zheng et al., 2022) uses it for its periphery, and FHL (Liu et al., 2021, 2022a) use it as a forest; 5) All Pair Pre-computation pre-computes all-pair distance with the fastest query, but incurs long index construction and large index storage. G-Tree (Zhong et al., 2013, 2015) and ROAD (Lee et al., 2009, 2010) store the distance between boundary vertices in each partition.
3. Novel PSP Strategies
As stated in Section 2, the traditional PSP strategy can be time-consuming for large networks since it requires a laborious computation of all-pair boundary distances using Dijkstra’s search. To tackle this problem, we propose two novel PSP strategies and analyze how they could be leveraged to correctly answer the shortest path queries. Besides, we design the index update approaches systematically to support the dynamic application of both the existing and novel PSP strategies. Lastly, we proposed a pruning technique to slim the overlay graph to achieve better index performance. It should be noted that the PSP strategy is a general procedure that any PSP index needs to follow and can work with any type of graph partition method or SP algorithm. So we introduce the PSP strategies without mentioning the specific partitioning and SP approach.
3.1. No-Boundary and Post-Boundary Strategy
Since the traditional partitioned index starts by precomputing the all-pair distance among boundary vertices, we call this approach Pre-Boundary Strategy. It seems that this strategy, with the theoretical guarantee for correct partitioned shortest-path processing, could handle all the partitioned queries well. Nevertheless, the first step of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized. As a result, the index construction and maintenance efficiency suffer when the graph has numerous boundary vertex pairs. However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index, otherwise the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition. For instance, as shown in Figure 3, $p(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$. As a result, $d(v_{6},v_{11})$ cannot be answered correctly only with $G_{3}$ and the global shortest distance $d(v_{6},v_{11})$ calculation is essential before constructing $L_{3}$. Then a question naturally arises: do we really require precomputing the boundary all-pair distance first? In other words, does there exist a chance to keep index correctness by dropping this time-consuming step? Based on this brainstorm, we break the traditional misconception and propose a novel No-Boundary Strategy as follows, which significantly reduces the index construction/maintenance time by skipping the time-consuming pre-computation step.
Index Construction. It contains three steps as shown in Figure 4. With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the shortest path index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$. For instance, $\tilde{G}$ is made up of boundary all-pair edges $(v_{3},v_{10})$, $(v_{8},v_{9})$, $(v_{5},v_{12})$, $(v_{6},v_{11})$ (with their weight calculated based on $L_{1},L_{2},L_{3}$, and $L_{4}$) and inter-edges. It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the shortest path index $\tilde{L}$ for $\tilde{G}$. Since the partition indexes $\{L_{i}\}$ are constructed in parallel first followed by the construction of $\tilde{L}$, the No-Boundary takes $max\{O_{c}(G_{i})\}+O_{c}(\tilde{G})$ time in index construction.
Figure 4. No-Boundary and Post-Boundary Strategies
Query Processing. Now that $L_{i}$ cannot answer $G_{i}$’s query correctly as the global distance between boundary all-pair is not covered, then how can No-Boundary answer query correctly? Before revealing it, we first prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, that is the queries between boundary vertices can be correctly processed with $\tilde{L}$, using:
Theorem 1.
$\forall s,t\in B,d_{G}(s,t)=d_{\tilde{L}}(s,t)$.
Proof.
We divide all the scenarios into three cases as shown in Figure 5-(a): 1) $s,t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds. Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s,t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$. In the latter two cases, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). For two adjacent vertices $b_{i},b_{j}\in p_{c}$, if $b_{i}$ and $b_{j}$ are in the same partition, then $\tilde{w}(b_{i},b_{j})$ can be correctly obtained in $\tilde{G}$ via case 1); otherwise, $(b_{i},b_{j})$ is an inter-edge with $\tilde{w}(b_{i},b_{j})=$$w(b_{i},b_{j})$ naturally correct. Thus, the shortest distance among boundary vertices can be correctly calculated by accumulating the edge weights on $\tilde{G}$. ∎
Figure 5. Different Categories of OD Distribution
Based on the above Theorem, we continue to answer the shortest distance queries by discussing all scenarios in two cases:
Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ as follows:
Theorem 2.
[…]
Proof.
We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and divide all the scenarios into two subcases:
Subcase 1: $p_{G}(s,t)$ does not go outside of $G_{i}$, as shown in the first row of Figure 5. No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path.
Subcase 2: $p_{G}(s,t)$ passes outside of $G_{i}$, as shown in the second row of Figure 5. If $s$ and $t$ are both boundary vertices, $d_{G}(s,t)=d_{\tilde{L}}(s,t)$ holds by referring to Theorem 1. If $s$ and $t$ are both non-boundary vertices, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). Therefore, $d_{4}$ can correctly handle this case as the $d_{G}(b_{0},b_{n})$ can be answered by $d_{\tilde{L}}(b_{0},b_{n})$ as per Theorem 1, while $d_{G}(s,b_{0})$ and $d_{G}(b_{n},t)$ can be answered by $d_{L_{i}}(s,b_{0})$ and $d_{L_{i}}(b_{n},t)$ by referring to Case 1. If either $s$ or $t$ is a non-boundary vertex, its distance is the special case of $d_{4}$ and can be easily proved. ∎
Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process queries the same manner as that of Pre-boundary Strategy.
Lemma 3.
The cross-partition queries can be correctly processed in the No-boundary Strategy.
Note that the proof of Lemma 3 and all the other Lemmas of this paper are available in the Appendix of our extended version (Zhang et al., 2023b). We can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The intra-query complexity is $max\{O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$, while the inter-query takes $O_{q}(\tilde{L})$ time when $s$ and $t$ are both boundary vertices, and the complexity of other case is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. Therefore, we make a bold attempt to enhance the performance of PSP index by skipping the heavy pre-computation, such that the index construction time is largely reduced so as the index maintenance (as will introduced in Section 3.2). Surprisingly, with incorrect distance value in the No-boundary index, we prove theoretically that it can support the query answering with correctness guarantee.
By comparing the query processing procedure of the traditional Pre-boundary strategy and our proposed No-boundary strategy, we find that it is more complex for No-boundary to process those queries with endpoints in the same partition because of the incorrectness of $\{L_{i}\}$. We repair this weakness by fixing the incorrect boundary all-pair distance in each partition $G_{i}$ and then transforming $\{L_{i}\}$ to their correct version. Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$. For instance, $d(v_{3},v_{10})$, $d(v_{8},v_{9})$, $d(v_{5},v_{12})$ and $d(v_{6},v_{11})$ are calculated based on $\tilde{L}$ and these edges (as colored red in Figure 4) are inserted into their corresponding partitions with new subgraphs $\{G^{\prime}_{i}\}$ formed. Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022) in each partition parallelly. As these procedures are implemented after the No-boundary, we call this strategy as Post-boundary. By referring to Theorem 1, the correct global distance information is contained in $\{G^{\prime}_{i}\}$, so $\{L^{\prime}_{i}\}$ can support the correct query processing of $Q(s,t)$ with $s,t\in G_{i}$. With the consistent query procedure as that of Pre-boundary, the query efficiency of Post-boundary reaches the state-of-the-art level.
3.2. Index Update for Different PSP Strategies
The partitioned shortest path has been widely used, however, how to maintain the index to support its application in dynamic environments has never been discussed. In this section, we propose the partitioned index update algorithms to guarantee the correctness of the partitioned index in dynamic scenarios.
Index Update for Pre-boundary Strategy. As shown in Figure 6-(a), when the weight of edge $e\in G$ changes, we first recalculate the boundary-all pair distance (as the step 1 in index construction) and identify the changed weight $e(b_{i1},b_{i2})$ between boundary vertices in each partition $G_{i}$. Then we need to update the corresponding partition index $L_{i}$ and overlay index $\tilde{L}$.
Lemma 4.
The indexes of Pre-Boundary Strategy can be correctly updated with the above strategy.
Step 1 takes $O_{u}(nlogn+m)$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
Figure 6. Index Update of Different PSP Strategies
Index Update for No-boundary Strategy. Since the weight change of inter-edges does not affect the index for each subgraph, we divide index updates into two scenarios as shown in Figure 6-(b). Scenario 1: Inter-edge weight change. When $e\in E_{inter}$ changes, only $\tilde{L}$ needs an update; scenario 2: Intra-edge weight change. When $e\in E_{intra}$ ($e\in E_{j}$) changes, we first update $L_{j}$ and compare the old and new weights of $e(b_{j1},b_{j2})$ between boundary in $G_{j}$. If there is an edge weight update, we need to further update $\tilde{L}$. The update complexity is $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})$.
Lemma 5.
The indexes of No-Boundary can be correctly maintained with the above update strategy.
Index Update for Post-boundary Strategy. It is similar to No-Boundary, with an additional judgment and processing as shown in Figure 6-(b). Scenario 1: Intra-edge weight change. Suppose $e\in E_{i}$ changes, we update $G_{i},L_{i}$ and then update $\tilde{G},\tilde{L}$ if any $d_{L_{j}}(b_{j1},b_{j2})$ changes. After that we may need to further update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; scenario 2: Inter-edge weight change. Suppose $e\in E_{inter}$ changes, we update $\tilde{G},\tilde{L}$ and then update $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$. Therefore, the update time complexity is the linear combination of the above procedures: $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})+O_{q}(\tilde{G})+max\{O_{u}(G_{i})\}$.
Lemma 6.
The indexes of Post-Boundary Strategy can be correctly maintained with the above update strategy.
Remark 1. To construct the PSP index, Pre-Boundary precomputes the all-pair shortest distance among boundary vertices of each partition, and thus the newly added edges of $G_{i}$ and $\tilde{G}$ (red edges) are of correct edge weight. For example, the edge weight of $e(v_{5},v_{12})$ is 6 which is the path length of the shortest path $p_{G}(v_{5},v_{12})=\left<v_{5},v_{3},v_{10},v_{12}\right>$. While No-Boundary skips the heavy precomputation and only leverage $L_{i}$ to calculate boundary all-pair distance for overlay graph $\tilde{G}$ construction. As a result, the edge weight $w_{\tilde{G}}(v_{5},v_{12})=7$ is incorrect since it only preserves the local distance with corresponding lcoal shortest path $p(v_{5},v_{12})=\left<v_{5},v_{4},v_{12}\right>$ in $G_{3}$. This innovation largely saves the index construction time while drags down the query answering speed. Then Post-Boundary makes up the query efficiency loss by inserting the correct shortest distance value between boundary pairs into each partition (e.g., insert $w(v_{5},v_{12})=6$ to $G^{\prime}_{3}$) and refresh $\{L_{i}\}$ to its correct version $\{L^{\prime}_{i}\}$. In summary, there exists a balance of index performance among the three boundary strategies. In terms of the query processing, the No-Boundary is slower than Pre-Boundary and Post-Boundary, since the boundary vertices and path concatenation should be considered for OD within one partition. In terms of the index construction and update, the No-Boundary and Post-Boundary enjoy faster speed as the boundary all-pair distance computation can be neglected. In terms of the index size, the Post-Boundary needs twice the storage space as that of No-Boundary and Pre-Boundary, even though it has advantages in both query processing and index update. Therefore, no one strategy is better than the others in all aspects and could be selected based on the application scenario. Besides, it is worth noting that we illustrate PSP strategies in the form of edge-cut for clear presentation since the vertex-cut partitions can be easily converted to edge-cut partitions. Due to limited space, the conversion method is presented in extended version (Zhang et al., 2023b).
3.3. Pruning-based Overlay Graph Optimization
Though it is flexible to select the suitable boundary strategy according to application scenarios, we observe that all boundary strategies are hindered by the extremely dense overlay graph. The density of overlay graph $\tilde{G}$ increases dramatically as the all-pair boundary vertices in each partition are connected during its construction, which badly affects the index performance by slowing down the index construction, query processing, and index update (Zhang et al., 2021a). Then is it possible to improve index performance and decrease the density of $\tilde{G}$ by deleting some unnecessary edges in $\{B_{i}\times B_{i}\}$? As proved in Theorem 1, the shortest distance between boundary vertices is well-preserved in $\tilde{G}$, then could the distance still be preserved in $\tilde{G}$ if some edges are removed? We start exploring this question by categorizing the boundary vertices as followed:
Definition 0 (Half/Full-Connected Boundary Vertex).
$b\in B_{i}$ is a half-connected boundary vertex if $\exists u\in N(b),u\in G_{i},u\notin B_{i}$ and we denote them as $B_{i}^{H}$ with $B^{H}$=$\{B_{i}^{H}\}$; otherwise, $b$ is a full-connected boundary vertex and denoted as $B_{i}^{F}$ with $B^{F}$=$\{B_{i}^{F}\}$.
Figure 7. Example of Overlay Graph Simplification
As illustrated in Figure 7-(a), $b_{i3}$ is a half-connected boundary vertex since there exists $v_{i1}\in N(b_{i3}),v_{i1}\notin B_{i}$; and $b_{i1},b_{i2},b_{i4}$ are full-connected boundary vertices. Initially, $\tilde{G}$ can be viewed as composed of multiple complete graphs $\{B_{i}\times B_{i}\}$, which are connected by the inter-edges. Within this dense structure, the global distance between any two boundary vertices $s,t\in B$ is well-preserved. Let us consider a shortest path $p$ with starting vertex $b_{s}\in B_{i}^{F}$ and target vertex $b_{t}\in B$, then the second vertex $v_{0}$ along $p$ is a boundary vertex with $v_{0}\in B,v_{0}\in N(b_{s})$ according to the definition above. Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{s},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{s})$, since it is impossible for $p$ to pass through these edges. The same principle holds for a shortest path ending with a fully connected boundary vertex $b_{t}\in B_{i}^{F}$ that its distance will also be preserved in $\tilde{G}$ after deleting edges $\{(b_{i},b_{t})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{t})$. Then for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph. As analyzed above, we can shrink $\tilde{G}$ ($\{B_{i}\times B_{i}\}\cup E_{inter}$) to its slimmer version $\tilde{G}^{\prime}$ (composed of $\{B_{i}^{H}\times B_{i}^{H}\}$, $E_{inter}$ and $B_{i}^{F}$’s adjacent edges). Suppose $b_{i1},b_{i2}\in N(b_{i4})$ and let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 7-(b) will be shrunk to a sparser one in Figure 7-(c) after the pruning technique. The following lemma demonstrates that the global distance can still be preserved in the overlay graph with the pruning technique.
Lemma 8.
$d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$, $\forall(b_{s},b_{t})\in B\times B$.
Therefore, we can enhance the index performance by pruning some edges in $\tilde{G}$ and utilizing its sparser version $\tilde{G}^{\prime}$, while preserving the global distance by referring to Lemma 8.
4. SP-Oriented Partition Method Classification
Having figured out how to achieve efficient query and update in partitioned graphs while guaranteeing the index correctness with the novel PSP strategies, we move on to the remaining two dimensions. As the SP index is well studied and analyzed in Section 2.4. Then, only the mist on the partition method remains: how to choose a partition method to achieve our preferred performance? This is not trivial because: 1) dozens of partition methods with different characteristics were proposed and applied in the past decades; 2) they are classified under different criteria (Buluç et al., 2016), for example, from the perspectives of partition manners (spectral, flow, graph glowing, contraction and multi-level), partition objectives (balance and minimal cut), computation manner (in-memory, distributed and streaming), and cut category (edge-cut and vertex-cut), but it is not clear which criterion is beneficial to SP index; 3) their relationship with (or the influence on) SP index is unknown and has never been studied. Based on these doubts, we propose a novel SP-oriented classification of partition methods as follows:
1) Planar Partition treats partitions equally on one level. It decomposes a graph $G$ into multiple equally-important subgraphs $\{G_{i}|1\leq i\leq k\}(k\in[2,\infty])$ with 1) $\bigcup_{i\in[1,k]}V(G_{i})=V$, $V(G_{i})\cap V(G_{j})=\emptyset$, or 2) $\bigcup_{i\in[1,k]}E(G_{i})=E$, $E(G_{i})\cap E(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$. Representative methods include spectral partitioning (Pothen et al., 1990; Barnard and Simon, 1994), graph growing (Diekmann et al., 2000), flow-based partitioning (Sanders and Schulz, 2011; Delling et al., 2011), node-swapping (Kernighan and Lin, 1970; Fiduccia and Mattheyses, 1982), multilevel graph partition (MGP) (Karypis and Kumar, 1998; Barat et al., 2018), etc..
2) Core-Periphery Partition ((Borgatti and Everett, 2000; Rombach et al., 2014; Elliott et al., 2020; Maehara et al., 2014a)) treats the partitions discriminately by taking some important vertices as “Core” and the remaining ones as “Peripheries”. It decomposes a graph $G$ into two distinct parts: a core subgraph $G_{c}$ and $k-1$ peripheral subgraphs $\{G_{i}|1\leq i\leq k-1\}(k\in[2,\infty])$, satisfying $\bigcup_{i\in[1,k-1]}V(G_{i})\cup V(G_{c})=V,V(G_{c})\cap V(G_{i})=\emptyset(\forall 1\leq i\leq k-1)$ and $V(G_{i})\cap V(G_{j})=\emptyset(\forall i,j\in[1,k-1],i\neq j)$. There are two big streams depending on how the core is formed: Core-Tree Decomposition (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022) forms the core through tree decomposition (Robertson and Seymour, 1984), and the resulting periphery part is a set of small-width trees. Specifically, it leverages minimum degree elimination (Xu et al., 2005) to contract vertices of lower degree first such that the graph is contracted from the edge towards the center, generating a set of growing trees (“Peripheries”) around a shrinking while denser graph (“Core”) formed by those non-contracted vertices. The contraction terminates when the width of one tree reaches the previously set threshold; Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) selects a set of vertices as landmarks, which can be regarded as the core and treats the remaining parts as periphery. It generally works on extremely large graphs where the above tree decomposition is impossible.
3) Hierarchical Partition organizes the network partitions hierarchically and each level is equivalent to planar partitioning. It organizes a graph $G$ hierarchically within $H$ levels where each level $h$ is a planar partitioning $\{G^{h}_{i}\}$ of $G$ and $\exists G^{h-1}_{j}\supset G^{h}_{i},\forall G^{h}_{i}(1\leq h<H)$. Classic methods include HiTi (Jung and Pramanik, 1996), SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a), SMOG (Holzer et al., 2009), and so on.A typical approach to create hierarchical partitions is by utilizing MGP, such as METIS (Karypis and Kumar, 1998) in G-Tree and G*-Tree.
Figure 8. Different SP-Oriented Partitioning
Table 1. SP-Oriented Partition Methods Classification.
Partition Structure Partition Method Cut Category Partition Objectives Spatial-Aware
Category Representative Methods Edge Cut Vertex Cut Balance Minimal Cut
Planar Spectral Partitioning SP (Pothen et al., 1990), RSB (Barnard and Simon, 1994) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Graph Growing Bubble (Diekmann et al., 2000) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Flow-Based Partitioning PUNCH (Delling et al., 2011) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Geometric Partitioning RCB (Simon, 1991; Huang et al., 1996) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Node-Swapping Heuristic KL (Kernighan and Lin, 1970), FM (Fiduccia and Mattheyses, 1982) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Minimum $k$-cut $k$-cut (Goldschmidt and Hochbaum, 1994) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Multilevel Graph Partitioning SCOTCH (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaHyPar (Akhremtsev et al., 2017) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Streaming Edge Partitioning HDRF (Petroni et al., 2015; Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
In-Memory Edge Partitioning NE (Zhang et al., 2017), HEP (Mayer and Jacobsen, 2021) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Core-Periphery Core-Tree Decomposition Core-Tree (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022; Elliott et al., 2020) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Sketch Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Hierarchical Edge-Cut Based HP HiTi (Jung and Pramanik, 1996) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Vertex-Separator Based HP SMOG (Holzer et al., 2009), ROAD  (Lee et al., 2009, 2010) $\times$ $\checkmark$ $\times$ $\times$ $\times$
Figure 8 shows the partition results of an example graph $G$ with different partitioning methods, where the red vertices represent the corresponding boundary vertices. Specifically, planar partitioning generates four partitions with equivalent vertex size while core-periphery partitioning produces both “Core” and “Periphery” (the periphery of Core-Tree decomposition is a set of small-width trees while Sketch treats the non-core part as periphery. The yellow vertex in each periphery is the “root vertex”). Hierarchical partitioning (HP) organizes partitions hierarchically, and the leaf nodes (the lowest level partition result) may have the same partition result as planar partitioning if the same partitioning method is used. In addition, we summarize and categorize existing partition methods under our SP-oriented classification, as shown in Table 1.
Remark 2. As analyzed in Section 3, the boundary number of each partition will affect the partition method section as it is crucial to the computation complexity of PSP index. Specifically, the planar and hierarchical partitions have no limit to the boundary number, even though reducing the cut size is one of their optimization goals as the path index may suffer from large boundary numbers. Complementary, the core-tree decomposition limits the boundary number to be no larger than the pre-defined bandwidth such that their performance would not be deteriorated by the boundary if we set the bandwidth wisely. Therefore, the planar and hierarchical partitions are better used for small treewidth networks such that balanced partitions and fewer boundaries can be achieved together; while core-periphery is suitable for large treewidth networks as it deliberately limits the boundary vertex number by bandwidth.
5. Universal PSP Index Scheme and Newly Generated PSP Indexes
In this section, we propose a universal PSP index scheme and generator and put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
Figure 9. PSP Index Generator with PSP Index Scheme and Representative Generated PSP Indexes
5.1. PSP Index Scheme and Generator
After elaborating on each component of the PSP index, we now have our PSP index scheme ready as shown in Figure 9. The main part is the PSP index generator, which is formed by coupling those three components. For easy selection of the PSP strategy and SP algorithm, we sort all of their choices roughly in terms of query and update efficiency. As shown in Figure 9, the lighter color signifies faster efficiency for PSP strategy and SP algorithm. Meanwhile, planar and hierarchical partitions are more suitable for low-treewidth networks, while core-periphery can deal with high-treewidth graphs. Given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator. For starters, the existing PSP indexes can find their positions in our scheme as will be discussed in Section 7, and we can compare them theoretically and fairly. Apart from this, we can further utilize it as a PSP Index Generator to create new PSP indexes that are suitable to user requirements, especially when all the existing ones are query-oriented, while there leaves a vast space of combinations for new ones. Specifically, the generated PSP index needs to choose the components from each of the three dimensions, and its name also has three parts corresponding to the three dimensions. Although we can enumerate them all, we choose to introduce the representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9.
5.2. Representative Generated New PSP Indexes
1. Query-Oriented Planar PSP Index: P-TD-P. Given a planar partition, the PSP strategies can apply directly to it. Next, we introduce how to construct the PSP indexes that are efficient in query processing under planar partition. Firstly in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is faster than Pre-Boundary in index construction, we use it as the PSP strategy. Secondly, in terms of the SP algorithm, we could choose TD as the index for both partitions and overlay graph. Although all-pair is faster, its space consumption is intolerable. Its structure is shown in Figure 9-(a). Specifically, the index construction takes $O(V_{max}\cdot(\log V_{max}+h_{max}\cdot w_{max}))$ time for the partition TD, $O(V_{\tilde{G}}(\log V_{\tilde{G}}+h_{\tilde{G}}\cdot w_{\tilde{G}}))$ for the overlay TD, $O(B^{2}_{max}w_{\tilde{G}})$ for boundary correction, and $O(w_{\tilde{G}}+w^{2}_{max}\cdot\delta)$ for the partition TD refresh, where $\delta$ is the affected shortcut number and $max$ are the corresponding max values in the partitions. For the query processing, the intra-query takes $O(w_{max})$ as partition index is correct, and the inter-query takes $O(max\{B_{max}w_{max},B^{2}{max}w_{\tilde{G}}\})$ for the partition and overlay query and $O(B^{2}_{max})$ for the combinations. For the index update, it takes $O(w^{2}_{max}\delta)$ to update partition TD, $O(w^{2}_{\tilde{G}}\delta)$ to update overlay TD, and $O(B^{2}_{max}w_{\tilde{G}})$ to check boundaries.
2. Update-Oriented Planar PSP Index: N-CH-P. In terms of the PSP strategy, we use No-Boundary as it requires the least effort to update. In terms of SP algorithm, we can choose CH as the underlying index as it is fast to update while the query processing is better than direct search (Figure 9-(b)). Its index construction is faster with $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for partition CH and $O(V_{\tilde{G}}\cdot w^{2}_{\tilde{G}}\cdot\log V_{\tilde{G}})$ for overlay CH; the query time is longer with $O(max\{B_{max}\cdot w_{max}\log V_{max},$ $B^{2}_{max}\cdot w_{\tilde{G}}\log V_{\tilde{G}}\})$ for the intra- and overlay searching, and $O(B^{2}_{max})$ for the combinations; the index update is faster with $O(\delta w_{max})$ for partition CH maintenance and $O(\delta w_{\tilde{G}})$ for overlay CH maintenance.
3. Query-Oriented Core-Tree PSP Index: P-PT-CP. The core-periphery partition index comprises the core index $L_{c}$ and the periphery index $\{L_{i}\}$. It seems that $L_{c}$ and $L_{i}$ can be constructed by their corresponding subgraphs $G_{c}$ and $G_{i}$, respectively. Although the core does not belong to any partition, they are connected to the partitions, and we treat the core as the overlay graph. Different from the previous planar partition, the core here usually has a large degree so its index is limited to PLL. As for sketch, we omit it here because its PLL core + pruned direct search seems to be the only solution for huge networks. Next, we discuss the remaining parts. For the PSP strategy, Post-Boundary is utilized as the queries within the periphery can be handled without the core index. The periphery index uses TD because the periphery usually has a small degree. This structure is shown in Figure 9-(c). In terms of index construction, because the periphery is constructed through contraction, we regard it as a by-product of the partition phase and do not construct their labels in the first step. Then it takes $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ for the core PLL, $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction, and $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for periphery label. In terms of query processing, the intra-query takes $O(w_{max})$ time. The inter-query takes $O(max\{B_{max}w_{max},B^{2}_{max}w_{c}\log V_{c}\})$ for the periphery and core, and $O(B^{2}_{max})$ for the combinations. In terms of index update, it takes $O(w^{2}_{max}\cdot\delta)$ for the periphery TD, $O(w_{c}E_{c}\log V_{c})$ for the core PLL, and $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction.
4. Update-Oriented Core-Tree PSP Index: N-PC-CP. As shown in Figure 9-(d), N-PC-CP uses CH for faster periphery update while PLL is for the core. It adopts No-Boundary strategy for faster update. In index construction, only core needs $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ time. The query time is longer with $O(max\{B_{max}\allowbreak w_{max}\log V_{max},$ $B^{2}_{max}w_{c}\log V_{c}\})$ for intra and core, and $O(B^{2}_{max})$ for combinations. The index update is fast with $O(\delta w_{max})$ for periphery shortcuts and $O(w_{c}E_{c}\log V_{c})$ for the core.
5. Update / Query-Oriented Hierarchical PSP Index: N-TS-HP. This category organizes $L$ levels of partitions hierarchically with several lower partitions forming a larger partition on the higher level. We use $L^{l}_{i}$ to denote the index of partition $G^{l}_{i}$ on level $l$. Different from the SOTA G-Tree stream of indexes which uses all-pair in their hierarchical overlay graph, we replace it with the hierarchical labels for better query and update performance. Specifically, for vertices in each layer, we store their distance to vertices in their upper layers. As this is essentially 2-hop labeling, we use TD to implement it with orderings corresponding to the boundary vertex hierarchy. Such a replacement in the overlay index could answer queries and update much faster than the original dynamic programming-based layer all-pair index. As for the partitions, this structure tends to generate small partitions so we inherit the original search for fast query processing. Consequently, no partition index leads to inevitable boundary all-pair searches. Fortunately, our No-Boundary restricts the search space to the small partition compared with G-Tree’s whole graph Pre-Boundary (Figure 9-(e) and (f)). To construct the index, the partition boundary all-pair takes $O(B_{max}\cdot V_{max}(logV_{max}+E_{max}))$, and its by-product boundary-to-partition can be cached for faster inter-query. Then the overlay TD takes $O(V_{\tilde{G}}(logV_{\tilde{G}}+h_{\tilde{G}}w_{\tilde{G}}))$ time. In query processing, the intra-query takes $O(B_{max}V_{max}(logV_{max}+E_{max})$ for the direct search (very rare as the partitions are small), while the inter-query The query takes $O(B_{max}w_{\tilde{G}}\})$ for intra- (constant time with cache) and hierarchical query, and $O(B^{2}_{max})$ for combination. As for index update, it takes $O(B_{max}\cdot V_{max}(logV_{max}+E_{max}))$ to update the partition all-pairs and $O(w^{2}_{\tilde{G}}\cdot\delta)$ to update the overlay graph.
Remark 3. We note that these five PSP indexes are novel structures, each with unique technical challenges, especially in updating. Nevertheless, we believe the PSP scheme holds greater significance than these new indexes, as it can guide new PSP index designs. We omit their details due to limited space, but their source codes are provided (GitHub, [n.d.]). Besides, we take the N-CH-P index as an example and elaborate on its details in the Appendix (FullVersion).
Table 2. Dataset Description
[] Type Name Dataset $|V|$ $|E|$ Road Network NY 1 New York City 264,346 730,100 FL 1 Florida 1,070,376 2,687,902 W 1 Western USA 6,262,104 15,119,284 US 1 Full USA 23,947,347 57,708,624 Complex Network GO 2 Google 855,802 8,582,704 SK 3 Skitter 1,689,805 21,987,076 WI 3 Wiki-pedia 3,333,272 200,923,676 FR 2 com-Friendster 65,608,366 3,612,134,270 [1] http://www.dis.uniroma1.it/challenge9/download.shtml [2] http://snap.stanford.edu/data [3] http://konect.cc/
Table 3. Average Vertex Degree of Overlay Graph
Dataset NY FL GO SK
Method PUNCH HEP PUNCH HEP KaHyPar HEP KaHyPar HEP
w/o Opt 48.68 325.98 36.17 719.25 947.36 9754.52 1420.18 34889.5
w/ Opt 48.68 320.61 36.17 714.30 912.24 7678.05 1409.98 27326.5
6. Experimental Evaluation
In this section, we evaluate the effectiveness of the proposed methods and provide a systematic evaluation of PSP indexes. All the algorithms are implemented in C++ with full optimization on a server with 4 Xeon Gold 6248 2.6GHz CPUs (total 80 cores / 160 threads) and 1.5TB memory. The thread number is set to 150.
6.1. Experimental Settings
Datasets and Queries. We test on eight real-life datasets, including four weighted road networks (NY, FL, W, and US) and four complex networks (Web graph: GO and WI; Social network: SK and FR), as shown in Table 2. The edge weight of complex networks is generated by following the manner (Zhang et al., 2021b) that is inversely proportional to the highest degree of the endpoints. We randomly generate 10,000 queries and 1,000 update instances for each dataset to assess the query processing and index maintenance efficiency, respectively.
Partition Methods and Their Performance Metrics. To select the suitable partition method for PSP index, we implement 8 representative partition methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar partition methods with PUNCH and RCB only working on road networks. Since core-periphery partition (including Tree-Decomposition-based partition (Ouyang et al., 2018) and landmark selection in Sketch (Farhan et al., 2022)) and hierarchical partition (mostly using METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index. We evaluate the performance of a partition method from three aspects: overall border number $|B|=\sum_{i\in[1,k]}|B_{i}|$, average partition border number $\overline{|B_{i}|}=\frac{\sum_{i\in[1,k]}|B_{i}|}{k}$, and partition connectivity $R_{C}=\frac{\sum_{i\in[1,k]}|C(G_{i})|}{k}$, where $C(G_{i})$ is the connected component number of subgraph $G_{i}$.
PSP Indexes and Their Performance Metrics In terms of PSP indexes comparison, we evaluate 5 existing representative PSP indexes spreading over our proposed three partition structures: a) FHL (Liu et al., 2021, 2022a) and THop (Li et al., 2019c) belongs to planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) and Sketch (Farhan et al., 2022) belong to the core-periphery partition; c) G-Tree (Zhong et al., 2013) belongs to the hierarchical partition. Besides, we implement our 5 newly proposed PSP indexes (see Section 5): P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP. For the state-of-the-art shortest path algorithms CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL respectively) to test the performance of different PSP strategies and select the suitable partition number. We measure the performance of the PSP index from four aspects: index construction time $t_{c}$, query time $t_{q}$, update efficiency $t_{u}$, and index size $s$. Note that we do not report the result (or set it to $\infty$) if the running time of a partition method or PSP index exceeds 24 hours.
Parameter Setting. We set the default partition number as $32$ for all planar partitions and the bandwidth of core-tree indexes as $40$ as per experimental results. The landmark number of Sketch is set to $20$ by referring to (Farhan et al., 2022). As for the hierarchical partition, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the maximum leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US, respectively.
Figure 10. Partition Performance When Varying $k$.
Figure 11. Effect of Partition Methods. Bar: Left, Ball: Right
Figure 12. Effect of Partition Number. Bar: Left, Ball: Right
Figure 13. Comparison with Non-partitioned SP Indexes. Bar: Left, Ball: Right
Figure 14. Effectiveness of Partitioned Shortest Path Strategy. (Bar: Left Axis; Ball: Right Axis)
6.2. Experiment Result Analysis
Exp 1: Performance of Partition Methods When Varying $k$. We first report the performance of different partition methods with varying partition number $k$ (we only report results on FL, US, and GO due to limited space; the full results are in the full version (Zhang et al., 2023b))). As shown in Figure 10 (a)-(f), the overall border number $|B|$ generally increases, while the average border number $\overline{|B_{i}|}$ decreases as $k$ grows from 4 to 256. This indicates that having a larger partition number can potentially speed up the query processing of the PSP index due to a smaller average partition border number. However, it also leads to a larger overlay graph, which can slow down the query processing and construction of the overlay index. Therefore, there should be a trade-off between them, and the best partition number typically varies for different PSP indexes. Furthermore, as per Figure 10 (g)-(i), only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1, which implies that at least one partition generated by other partition methods is unconnected (with multiple separate components). Since PTD requires the vertices within one partition to be connected, only PUNCH and Bubble suit it.
Exp 2: Effect of Partition Method. Next, we further evaluate the effect of the partition method on the PSP indexes (PCH, PTD, and PPLL). In particular, we test PCH and PPLL with Bubble, PUNCH, KaHyPar, and METIS on road networks (NY and FL) while only testing PTD with PUNCH and Bubble by referring to Exp 1. Figure 11 shows that PUNCH and KaHyPar perform better than other partition methods in most cases due to smaller overall border number $|B|$ and average partition border number $\overline{B_{i}}$, which indicates that $|B|$ and $\overline{B_{i}}$ are important factors in the partition method selection. We select PUNCH as the default planar partition method due to its superior PSP index performance. It is worth noting that we also test PPLL with KaHyPar, METIS, and SCOTCH on complex networks. However, it fails to complete the index construction on the smallest graph GO within 48 hours due to the large and condensed overlay graph, which reveals that the PSP index with planar partition is generally unsuitable for complex networks.
Exp 3: Effect of Partition Number. As shown in Figure 12 (a)-(b), the indexing time of PCH and PTD increased as $k$ increased from 4 to 256. This is because a larger $k$ could lead to a larger overlay graph that cannot leverage thread parallelization, thus resulting in higher indexing time. By contrast, PPLL’s indexing time first decreases and then increases in NY while tending to decrease in FL because a larger $k$ could also lead to a smaller $\overline{|B_{i}|}$, thus reducing the overlay graph construction time. In Figure 12 (c)-(d), we observe that a smaller $k$ generally improves the query time but leads to worse index update efficiency. Therefore, it is crucial to balance the query and update efficiency by selecting the appropriate partition number. We set the default value of $k$ as 32 since it provided satisfactory performance in most cases.
Exp 4: Comparison with Non-partitioned SP Algorithms. We conduct a comparison between PSP indexes (including PCH, PTD, PPLL, and CT) and their non-partitioned counterparts (namely CH (Geisberger et al., 2008), TD (Ouyang et al., 2018), PLL (Akiba et al., 2013)). As depicted in Figure 13, PSP indexes generally provide faster indexing time, smaller index size, and better scalability than their non-partitioned counterparts. For instance, PCH and PTD exhibit up to 2.76×2.76$\times$2.76 × and 2.25×2.25$\times$2.25 × faster index construction compared to CH and TD, respectively. Moreover, CT demonstrates 72.3×72.3$\times$72.3 × and 3.58×3.58$\times$3.58 × improvement over PLL in terms of indexing time and index size, respectively. However, PSP indexes may compromise query processing efficiency and index update time. It is worth noting that it also shows that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as PPLL does not outperform PLL in terms of indexing time and scalability.
Exp 5: Effectiveness of PSP Strategy. As shown in Figure 14 (a)-(d), the no-boundary strategy has the minimum index construction (about $14-35\%$ speed-up) and index update time (up to two orders of magnitude speed-up compared with the pre-boundary strategy for PTD and PCH), which indicates the superiority of our no-boundary strategy. Suffering from the highest index construction time, the post-boundary strategy has the same query time as the pre-boundary strategy but a significant update efficiency improvement (14.636.7×14.6-36.7$\times$14.6 - 36.7 × speed up for PTD). To further demonstrate the effectiveness of our post-boundary strategy, we test the same-partition query efficiency as presented in Figure 14 (e)-(f). The post-boundary archives about 3.76.3×3.7-6.3$\times$3.7 - 6.3 × and 4.830.5×4.8-30.5$\times$4.8 - 30.5 × speed-up compared with the no-boundary strategy for PTD and PPLL, respectively. Such a significant performance gain comes from the post-boundary’s correct local partition index. Finally, we evaluate the effectiveness of pruning-based overlay graph optimization by measuring the average vertex degree of the overlay graph, as shown in Table 3. We test PUNCH and HEP on road networks while the KaHyPar and HEP on complex networks. The degree decreases with the pruning-based optimization in most cases. It cannot improve PUNCH since PUNCH already has a great partition result.
Figure 15. Comparison of All Shortest Path Indexes. Bar: Left, Ball: Right
Exp 6: Comparison of All PSP Algorithms. We compare our proposed PSP indexes with the existing PSP indexes in Figure 15. We first analyze the performance of our proposed indexes against their counterparts: 1) P-TD-P has the same query performance and index size as FHL, but it is much faster to update with $1$ 202.7 × speed-up. 2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL. 3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index. P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries. 4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction. 5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
We next discuss the advantages of different partition structures. Figure 15-(b) shows that the state-of-the-art planar PSP indexes (N-CH-P, P-TD-P) and hierarchical PSP indexes (N-TS-HP) generally provide better query and index update efficiency for road networks than core-periphery PSP indexes. However, only core-periphery PSP indexes can be used on complex networks, which typically have large treewidth. Therefore, planar and hierarchical partitions are more suitable for road networks, while complex networks work better with core-periphery partitions.
6.3. PSP Application Guidance
Based on the experimental analysis in Section 6.2, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) all PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm. Moreover, we could combine the elements of these three dimensions to create a new PSP index; 2) it is important to choose the right partition structure for different types of graphs. Generally, for road networks, planar and hierarchical PSP indexes are more suitable than core-periphery PSP due to their efficient query and index update efficiency. On the other hand, core-periphery PSP indexes are more suitable for complex networks due to their excellent scalability; 3) although PSP indexes typically have faster indexing efficiency, smaller index size, and better scalability than their non-partitioned counterparts, they sacrifice query and index update efficiency.
According to our theoretical analysis and experimental results, we further provide our decision tree for PSP index design as illustrated in Figure 16. The first step in creating a new PSP index is to select a partition structure that is suitable for the specific application, such as graph type and size. For graphs with low treewidth, such as road networks, we recommend using planar or hierarchical partition structures. For larger graphs, we suggest using the core-periphery structure, with the core-tree structure for million-level graphs and sketch for billion-level graphs. The second step is to select the PSP strategy based on the required query processing and index update efficiency. For query-oriented applications, we recommend using the post-boundary strategy, while for update-oriented applications, the no-boundary strategy is more suitable. The third step is to choose the SP algorithm for the overlay and partitions. For a small number of vertices, we recommend using the most efficient all-pair table. For larger graphs, if tree decomposition is feasible, we suggest using TD or CH to achieve better query/update efficiency. If tree decomposition is infeasible, we recommend selecting PLL or even the direct search for SP algorithm selection.
Figure 16. Our Decision Tree of PSP Index Design
7. Related Work
Recently, multiple surveys about the SP computation have emerged, focusing on either specific graph types (like transportation networks (Bast et al., 2016) or complex networks (Zhang et al., 2022)) or static graphs (Madkour et al., 2017; Li et al., 2017; Wu et al., 2012). (Zhang et al., 2021a) investigates the computation in dynamic graphs, but it does not consider the partitioned counterpart. The partitioned pathfinding framework is proposed in (Chondrogiannis and Gamper, 2014, 2016), however, the dynamic version is not discussed and they can be covered by the traditional PSP strategy. In this paper, we study the dynamic PSP problem systematically and explore its performance in both road networks and complex networks by involving the state-of-the-art discovery of SP algorithms and partition methods. Next, we discuss and categorize the existing PSP methods based on our scheme:
1) Pre-Boundary + Search/All-Pair + Hierarchy/Planar: These methods (HiTi (Jung and Pramanik, 2002), Graph Separators, Customizable Route Planning (Delling et al., 2017) (Holzer et al., 2009), ParDiSP (Chondrogiannis and Gamper, 2014, 2016)) pre-compute the shortest distance between boundaries (some hierarchically) to guide the search. In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category. G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching, and it is widely used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent routing (Wang et al., 2019), and machine learning-based path finding (Huang et al., 2021). They are slow to construct due to Pre-B and slower to query to direct search;
2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds the labels for the skyline shortest path on the overlay graph to answer the constrained shortest path query. This structure’s construction suffers from Pre-B and query suffers from searching;
3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022b, 2022a) partition builds the TD both within and between partitions for multi-dimensional skyline paths. As validated in Exp 4, our Post-B can speedup Pre-B’s construction dramatically;
4) Pre-Boundary+PLL/ PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long range time-dependent paths, and this structure’s performance is limited by PLL;
5) No-Boundary+ PLL/TD+Core: Core-Tree (Li et al., 2020a; Zheng et al., 2022) is the baseline that we already discussed;
6) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is nearly impossible so only a small number of landmarks are selected to either help prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate result (Gubichev et al., 2010; Qiao et al., 2012).
In sum, all the existing PSP methods have their limitation, warranting the comprehensiveinvestigation embarked on in this paper.
8. Conclusions
In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by concomitantly considering three dimensions: PSP strategy, SP algorithm, and partition structure. For the PSP strategy, we propose two new strategies and pruned-boundary optimization for better index performance. We also provide index maintenance solutions for both traditional and novel PSP strategies. For partition structure, we propose a new path-oriented partition method classification. Furthermore, we construct our PSP index scheme by seamlessly coupling those three dimensions. We also propose a PSP index generator and five new PSP indexes that are either more efficient in query or update than the current state-of-the-art. Systematic and extensive experimental evaluations fully validate our findings and prove the effectiveness of our new indexes guided by the insightful PSP index scheme. Finally, PSP index design guidelines are provided to help future researchers and practitioners design their indexes. Our future work will address recommendation aspects for such designs.